In a Series of Spring Training Games, the Yankees scored 35 runs. The Dodgers and Tigers, respectively, tallied half and a third as many as the Reds. The Reds' scores were as much below the Yankees' as the Dodgers' were above the Tigers'. How many runs did each team score?
Answer by mszlmb(115)

 Series of Spring Training Games, the Yankees scored 35 runs. The Dodgers
and Tigers, respectively, tallied half and a third as many as the Reds. The
Reds' scores were as much below the Yankees' as the Dodgers' were above the
Tigers'. How many runs did each team score? So Y=35 D=R/2 T=R/3 Y-R=D-T multiply equations 2 and 3 by 2 and 3, respectively. 2D=R 3T=R See that? 3T=R=2D, therefore, 3T=2D. So Y=35 3T=2D=R 35-R=D-T Let's mutate the second equation to our liking. R=3T=2D T=2/3D=R/3 D=R/2=1.5T Now, take D and T in the equation 35-R=D-T and replace them with their equivalents of R. 35-R=R/2-R/3 We're more than halfway done :D Now do R/2-R/3=3R/6-2R/6=(3R-2R)/6=R/6. Thence, 35-R=R/6. Now do +R on both sides. 35=R/6+R 35=7R/6 times 6 both sides 210=7R divide both by 7 30=R. The Reds tallied 30. Back to this: D=R/2 T=R/3 Replace R's with 30: D=30/2=15 T=30/3=10. The answers are: Yankees--35 Reds-----30 Dodgers--15 Tigers---10 Hope that helped!